whats the probability of getting a 2, 3, or 12 on two dice? whats the probability of not getting a 2, 3, or 12?
Answer by stanbon(48568)

whats the probability of getting a 2, 3, or 12 on two dice? whats the probability of not getting a 2, 3, or 12?
--------------
Draw a 6 by 6 square showing the sample space of the 36 possible
sums gotten from rolling two dice.
---------
Count the number of sums of 2: 1
Count the number of sums of 3: 2
Count the number of sums of 12: 1
--------------
P(2 or 3 or 12) = 4/36 = 1/9
Cheers,
stan H.